# UrbanVLP: A Multi-Granularity Vision-Language Pre-Trained Foundation Model for Urban Indicator Prediction

### v1

@@ Abstract.
Urban indicator prediction aims to infer socio-economic metrics in diverse urban landscapes using data-driven methods. However, prevalent pre-trained models, particularly those reliant on satellite imagery, face dual challenges. Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place. Secondly, the lack of interpretability in pre-trained models limits their utility in providing transparent evidence for urban planning. In response to these issues, we devise a novel Vision-Language Pre-Trained Model (UrbanVLP) in this paper. Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pre-trained models. Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery. Rigorous experiments conducted across six socio-economic tasks underscore UrbanVLP’s superior performance. We also deploy a web platform to verify its practicality.
Y. Liang is the corresponding author. Email: yuxliang@outlook.com
1. Introduction
Urban Indicator Prediction, a scholarly pursuit in the realm of urban computing (Zheng et al., 2014), deploys data-driven methodologies to forecast socio-economic metrics (e.g., GDP, population, carbon emission) in diverse urban landscapes. Rooted in the escalating global prominence of urban environments and the imperative for judicious urban planning, this academic discipline finds motivation in enhancing the efficacy of policymaking, optimizing resource allocation, and mitigating challenges endemic to urban agglomerations (Xi et al., 2022; Yan et al., 2023).
Classical machine learning approaches commonly tackle urban indicator prediction via supervised training, leveraging task-specific datasets like satellite imagery. These tasks span a spectrum from predicting poverty levels (Jean et al., 2016a; Han et al., 2020; Perez et al., 2017; Ayush et al., 2021a), crop yields (You et al., 2017; Wang et al., 2018; Martinez et al., 2021; Yeh et al., 2021), land cover (Uzkent et al., 2019), to commercial activity (He et al., 2018; Liu et al., 2023d). Nevertheless, the inherent task specificity of supervised learning, reliant on ample labeled data, may impede the model’s capacity for broad generalization. This constraint poses a conceivable risk to the overarching robustness and effectiveness of the model.
In contrast, Pre-trained Model, also known as Self-Supervised Learning (Jaiswal et al., 2020), empowers the autonomous discovery and optimization of urban region representations from diverse urban data modalities, encompassing satellite images (Jean et al., 2016b; Ayush et al., 2021b; Yao et al., 2022b; Watmough et al., 2019), Points of Interests (POIs) (Fu et al., 2019; Jenkins et al., 2019; Zhang et al., 2021; Li et al., [n. d.]), road networks (Liang et al., 2018, 2021; Wei Chen, 2023) and human mobility (Wang and Li, 2017; Seong et al., 2023; Li et al., 2023b, a). By obviating the necessity for externally labeled datasets, these pre-training techniques enable models to learn intrinsic region representations, fostering adaptability and scalability in comprehending intricate urban dynamics for enhanced generalization across various urban indicator prediction tasks.
Figure 1. Single-granularity vs. Multi-granularity modeling.
Given the widespread accessibility of satellite imagery through platforms such as Google Maps or Baidu Maps (Li et al., 2022b; Liu et al., 2023d; Xi et al., 2022), coupled with its wealth of information on regional features (e.g., road network patterns, building density, vegetation coverage), a predominant majority of initiatives harness satellite imagery [^1] As other modalities (e.g., mobility, POIs) are always inaccessible due to privacy issues. as the foundational modality for learning urban representations. Despite its promise, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture, storefront). As depicted in Figure 1, while the satellite images (a) and (b) possess remarkably similar layouts, their urban indicators considerably differ. Upon zooming into micro levels using the corresponding street-view images (c) and (d), a more nuanced and fine-grained understanding becomes apparent. Exclusively capturing macro-level patterns through satellite imagery may introduce substantial bias, falling short in furnishing more comprehensive information across diverse urban levels. Hence, the integration and alignment of multi-granularity information (macro-level and micro-level) for urban indicator prediction present an ongoing challenge that requires further exploration.
@@ 1. Introduction
Furthermore, while urban decision-makers rely on explainable evidence for service deployment and urban planning, the pre-trained models for urban indicator prediction mostly lack interpretability. They typically learn direct mapping from single or multi-modal data to the output representations, which poses a significant challenge in comprehending the intrinsic factors influencing the predictive outcomes. In light of the recent surge in Large Language Models (LLMs) (Ouyang et al., 2022; Zhao et al., 2023), UrbanCLIP (Yan et al., 2023) emerges as an innovative solution. This model effectively integrates textual modality as auxiliary knowledge into the realm of urban indicator prediction. Figure 1 shows how these textual data encapsulate key elements within both satellite images and their street-view counterparts. Notably, even in the absence of image-text pairs, UrbanCLIP adeptly harnesses a myriad of textual descriptions generated by pre-trained image-to-text LLMs, thereby facilitating both explainable reasoning and prediction. However, the text generation process of UrbanCLIP raises several critical concerns. a) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery. b) Homogenization: The generated description appears overly simplified and general, potentially leading to homogenization and inadequate differentiation. For instance, the average length of the text description in UrbanCLIP is quantified at 18.1 words, leading to insufficient information. To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
@@
In this paper, we present a Vision-Language Pre-Trained Model entitled UrbanVLP for urban indicator prediction. Our model elaborately integrates multi-granularity (macro and micro-level) information from both satellite and street-view imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions. Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module. It utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities. Compared to standard contrastive learning applied in the satellite branch at the macro level, we employ token-level contrastive learning to capture fine-grained details in street-view imagery. Concurrently, spatial information, enriched with the location coordinates of street-view images, is encoded and injected to provide essential spatial context. Finally, the fusion process aggregates features from street-view images and location-specific features within the same region, seamlessly incorporating fine-grained information into the satellite branch.
To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards. Given the impracticality of generating a large number of descriptions using GPT-4V (Yang et al., 2023) or Gemini (Fu et al., 2023) due to prohibitive expense, we utilize a closely comparable GPT4V (Yang et al., 2023) model called ShareGPT4V (Chen et al., 2023a), which is a strong caption model trained on 100k GPT4V (Yang et al., 2023) generated captions. To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric PerceptionScore, motivated by the human evaluation system. This metric comprises two components: Clipscore and Cyclescore, working collaborateively to capture both semantic similarity and detail inclusiveness.
Overall, the main contributions of our work are summarized as:
  • Multi-granularity cross-modal alignment. We explored the role of two different visual data modalities at different semantic granularities - street-view and satellite imagery. We introduced street-view data to inject fine-grained semantic information based on the satellite imagery and provided a framework.

### v2

# UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling

@@ Abstract
Urban region profiling aims to learn a low-dimensional representation of a given urban area while preserving its characteristics, such as demographics, infrastructure, and economic activities, for urban planning and development. However, prevalent pretrained models, particularly those reliant on satellite imagery, face dual challenges. Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place. Secondly, the lack of interpretability in pretrained models limits their utility in providing transparent evidence for urban planning. In response to these issues, we devise a novel framework entitled UrbanVLP based on Vision-Language Pretraining. Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pretrained models. Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery. Rigorous experiments conducted across six urban indicator prediction tasks underscore its superior performance.
1 Introduction
Urban Region Profiling, a scholarly endeavor within the domain of GeoAI [45, 54] and smart city [85], refers to the process of transforming urban regions into a low-dimensional vector representation using deep neural networks. Rooted in the escalating global prominence of urban environments and the imperative for urban planning, this academic discipline aims to capture and encode the multifaceted attributes of urban areas, such as demographic information, economic activity, environmental factors, infrastructure, and social dynamics, into a format that can be easily analyzed and utilized by computational models [73, 74]. Given the widespread accessibility of satellite imagery on platforms such as Google Maps [42, 51, 73], coupled with its wealth of information on regional features (e.g., road networks, building density, and vegetation coverage), a predominant majority of initiatives harness satellite imagery as the foundational modality for learning urban region representations [88].
Through a comprehensive review of the current literature, we summarize mainstream methods for urban region profiling into two categories: a) Task-specific supervised learning commonly realizes urban region profiling in a fully supervised task, e.g., identifying poverty levels [21, 59, 4, 5], crop yields [80, 70, 63, 57, 53, 78], and commercial activeness [24, 51]. However, its inherent task specificity (i.e., reliant on ample labeled data), may impede the model’s capacity for broad generalization to other downstream tasks. b) Vision pretraining, also known as a type of Self-Supervised Learning [28], aims to learn general visual features from urban satellite imagery, which are subsequently fine-tuned on a specific task for enhancing performance  [29, 6, 73, 51]. The inclusion of more auxiliary data, e.g., Points of Interests (POIs) [19, 29, 82, 40], road networks [46, 47, 12] and human mobility [71, 64, 44, 39], leads to richer, more accurate, and more useful representations of urban areas. For clarity, Figure 1(a-b) depict a sketch of these two streams of approaches.
Despite the success of using satellite imagery for urban region profiling, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture). Previous approaches in Figure 1(a-b) mainly focus on single granularity, neglecting finer-grained visual clues. Nonetheless, as the saying goes, “the devil is in the details.”. By zooming into micro levels using corresponding street-view images, a more nuanced and fine-grained understanding emerges, as shown in the left hand side of Figure 1(d). Therefore, the integration and alignment of multi-granularity information, both macro-level and micro-level, remain ongoing challenges in urban region profiling that require further exploration.
Figure 1: Frameworks for urban region profiling. Compared to existing literature, we present the first attempt to introduce multi-granular visual information and high-quality calibrated texts for this task.
@@ 1 Introduction
Furthermore, while urban planners rely on explainable evidence for service deployment and urban planning, the vision-pretraining-based models for urban region profiling mostly lack interpretability. They typically learn a mapping from single/multi-modal data to the representations, posing a significant challenge in comprehending the intrinsic factors influencing the prediction outcomes. In light of the recent surge in Large Language Models (LLMs) [58, 84, 55], UrbanCLIP [74] stands out as an innovative solution, which effectively integrates textual modality as auxiliary knowledge into urban region profile, as depicted in Figure 1(c). However, the text generation process of UrbanCLIP raises several critical concerns. i) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery. ii) Homogenization: The generated descriptions appear overly simplified and general, potentially leading to homogenization and inadequate differentiation. To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
@@
In this paper, we present a Vision-Language Pretraining framework (i.e., UrbanVLP) for urban region profiling. As depicted in Figure 1(d), our model elaborately integrates multi-granularity information from both satellite (macro-level) and street-view (micro-level) imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions. Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module, which utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities. To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards. Given the impracticality of generating a large number of descriptions using GPT4V [75] or Gemini [18] due to prohibitive expense, we utilize a closely comparable GPT4V [75] model called ShareGPT4V [11]. To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric called PerceptionScore, motivated by the human evaluation system.
In summary, our major contributions lie in the following aspects:
  • Multi-Granularity Cross-Modal Alignment. We explore the role of two distinct visual data modalities at various semantic granularities – satellite imagery and street-view imagery. We integrate street-view data to inject fine-grained semantic information through token-level contrastive learning. By aggregating and fusing local features within a specific region with satellite imagery data, we facilitate the development of a robust and comprehensive representation of urban landscapes.